OpenStudy (anonymous):

Please Help! It costs $350 to carpet an 8ft by 10ft room. At this rate, how many dollars would it cost to put the same carpeting in a 24ft by 30ft room? A. $380 B. $3,150 C. $1,050 D. $507 Ok, so I know that it's going to have something to do with ratios because that's what we are learning right now in this chapter. I can figure out that 8:10 is 4:5 and that 24:30 is 4:5, I can also figure out that 8:24 is 1:3 and that 10:30 is 1:3, only problem is that I have no clue what to do with this information :/ Am I on the right line of thinking? Help would be appreciated.

OpenStudy (anonymous):

@phi @Hero @mathmale @iPwnBunnies Help?

OpenStudy (anonymous):

@ganeshie8 Can you help?

OpenStudy (anonymous):

@thomaster ?

OpenStudy (anonymous):

350/80 = 35/8 dollars per square foot (35/8) (24)(30) = $3150

OpenStudy (anonymous):

Thanks!